# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v3

@@ IV.3 Forces and Stress
where the kinetic and nonlocal pseudopotential matrices are respectively labeled as $T_{\mu\nu}(\mathbf{R})$ and $V_{\mu\nu}^{\rm NL}(\mathbf{R})$, while the local pseudopotential is $V_{\rm local}$ and the ionic energy $E_{\rm II}$ is given by the Ewald method Martin (2020). It should be noted that ABACUS uses Eq. 33 to calculate the total energy in practical computations. For more information about evaluating force and stress terms in ABACUS, we refer the readers to a recent review work in Ref. Lin et al., 2024.
@@ IV.4 Hybrid Functional
Hybrid density functionals (HDFs) formulated within the generalized KS framework Seidl et al. (1996) belongs to the fourth rung of the Jacob’s ladder Perdew and Schmidt (2001). It overcomes the drawbacks of the local and semi-local XC functionals, namely the self-interaction errors Perdew and Zunger (1981); Mori-Sánchez, Cohen, and Yang (2008) related to the underestimation of the band gaps in solids. The ABACUS package supports pure exact exchange (EXX) calculations, specifically the Hartree-Fock (HF) method. Building upon this capability, we have implemented several widely-used hybrid functionals, including the Heyd-Scuseria-Ernzerhof (HSE) Heyd, Scuseria, and Ernzerhof (2003), PBE0  Ernzerhof and Scuseria (1999); Adamo and Barone (1999b), and SCAN0 Hui and Chai (2016). The local (semilocal) components of these functionals are computed using the LibXC library Lehtola et al. (2018), while the exact (Fock) exchange energy is evaluated through two-electron integrals of the Kohn-Sham orbitals.
@@
In fact, the above EXX effective potential exhibits an unfavorable $O(N^{4})$ scaling behavior, making it significantly more computationally demanding than local functionals in terms of both time and memory requirements. To mitigate this challenge, several algorithmic approaches have been developed, most notably the density fitting (DF) technique Whitten (1973) and its specialized implementation for localized atomic orbitals known as resolution of identity (RI) Feyereisen, Fitzgerald, and Komornicki (1993); Vahtras, Almlöf, and Feyereisen (1993); Ren et al. (2012). These methods reduce the four-center integrals to combinations of three- and two-center integrals by expanding the orbital products using auxiliary basis functions (ABFs).
For HDF calculations in large periodic systems, the local resolution of identity (LRI) approach Levchenko et al. (2015); Ihrig et al. (2015); Lin, Ren, and He (2020) enables linear-scaling EXX calculations by exploiting the inherent locality of both atomic orbitals (AOs) and ABFs. Extensive benchmarking has demonstrated that with properly chosen ABFs, the LRI approximation achieves sufficient accuracy for HDF calculations Lin, Ren, and He (2020). Residual discrepancies in cohesive properties and band gaps primarily originate from differences in core-valence interaction treatments (pseudopotentials versus all-electron methods) rather than basis set limitations Ji et al. (2022).
@@
Figure 20: Band gaps of 21 different semiconductors, calculated by ABACUS+LibRI, along with the reference computed values from other software (QE/FHI-AIMS/VASP) Lin, Ren, and He (2020) and experimental values Cui et al. (2018).
The computational efficiency is substantially enhanced by exploiting the inherent locality of both AOs and the density matrix, which enables effective prescreening of negligible submatrices below a specified threshold. For parallel computation of four-center integrals, we have implemented two complementary distribution strategies: a load-balancing greedy algorithm for multi-processor scheduling, and a spatially optimized scheme based on K-means clustering of ABF atom pairs. The latter minimizes both inter-processor communication overhead and memory requirements by ensuring geometrically proximate pairs are assigned to the same processor, with memory consumption scaling linearly with the union of their neighbor lists Lin, Ren, and He (2021a).
Notably, the LRI-based four-center integration has been modularized into the standalone LibRI library lib (2024a), which provides a portable and efficient implementation that can be readily integrated into various electronic structure codes. This includes both conventional DFT packages and beyond-DFT methods such as the random phase approximation (RPA) and GW calculations lib (2024b), offering substantial acceleration for two-electron Coulomb repulsion integral evaluations across different theoretical frameworks.
In 2025, Lin et al. developed a linear-scaling algorithm for exact-exchange forces and stresses in hybrid DFT, applicable to both molecules and periodic systems. Leveraging numerical atomic orbitals and localized resolution-of-identity, the massively parallel implementation enables structural relaxation of thousand-atom systems in ABACUS Lin et al. (2025).
@@
Generally speaking, hybrid functional calculations can only process systems of limited size due to the large computational demand of building the exact exchange Hamiltonian. By combining NAOs basis and LRI techniques, Fig. 20 shows the time consumption of this part increases almost linearly with the system size, enabling ABACUS to perform hybrid functional calculations for large systems with thousands of atoms. In 2024, Tang et al. used hybrid functional data from ABACUS and trained a deep equivariant neural network approach for efficient hybrid density functional calculations Tang et al. (2024). Taking twisted bilayer graphene (TBG) as an example in Fig. 21, they performed the HSE band structure calculations of (17, 16) TBG (twist angle $\theta\approx 2.0046^{\circ}$, 3,268 atoms/cell)Tang et al. (2024).
We also performed HSE calculations to obtain the band gaps of several semiconductors and compare them with the results from other DFT packages (QE/VASP/FHI-AIMS) and the experimental values Lin, Ren, and He (2020); Cui et al. (2018). In these tests, ABACUS and QE employ the SG15-type norm-conserving PBE pseudopotentials (except for In is PSlibrary norm-conserving), VASP employs the PBE PAW potentials, whereas FHI-AMIS performs all-electron calculations Delley (1990). Fig. 20 shows that the band gaps obtained by ABACUS HSE calculations are consistent with the results from other DFT packages, demonstrating satisfactory results when compared to the experimental values.
@@ IV.5 DeePKS
The exchange-correlation functional is critical to determining the accuracy of DFT. However, when different levels of XC functionals are selected, there exists a trade-off between accuracy and efficiency. Lower-level XC functionals on Jacob’s ladder Perdew (2001) typically offer higher computational efficiency but lower accuracy, whereas higher-level ones exhibit the opposite trend. With the rapid progress in increasing computational resources and advanced algorithms, AI-assisted methods hold the potential to mitigate this issue Kulik et al. (2022); Huang, Von Rudorff, and Von Lilienfeld (2023).
Proposed in 2020, DeeP Kohn-Sham method (DeePKS)Chen et al. (2021) employs a computationally efficient neural network-based functional model to represent the energy and force difference between a lower-level XC functional and a higher-level XC functional. The resulting model maintains translational, rotational, and permutational invariance and can be used in self-consistent field calculations. Together with DeePKS-kit Chen et al. (2022a) software, ABACUS supports the iterative training of DeePKS models Li et al. (2022a). Specifically, DeePKS models can be trained in molecular or periodic systems to learn properties such as energy, force, stress, and band gaps. Based on the DeePKS model, ABACUS can achieve accuracy similar to that of the high-level XC functional on these specified properties, while the computational efficiency is similar to that of the low-level XC functional.
@@
Here we list a few applications using the DeePKS method. For instance, halide perovskites (ABX3, X=halogen anion) have shown great promise as a cost-effective alternative to current commercial photovoltaic technologies. Designing effective photovoltaic systems requires a precise yet efficient description of the electronic structure of halide perovskites. One key benefit of halide perovskites is adjusting the absorption edge wavelength (band gap) by changing the ratio of different halide ions. Ou et al. Ou et al. (2023) constructed a general DeePKS model that can be utilized for halide perovskites, including different combinations of ABX3 (A=FA, MA, Cs; B=Sn, Pb; X=Cl, Br, I), the organic-inorganic hybrid alternatives, and the Ruddlesden-Popper (RP) perovskites. In detail, the authors built an extensive DeePKS model upon 460 configurations spanning seven types of halide perovskites, with HSE06 accuracy and satisfactory predictions for the band gap. Based on an iterative training process with ABACUS and DeePKS-kit, they showed that the resulting DeePKS model can accurately replicate forces, stress, band gaps, and density of states (DOS) near the Fermi energy for all types of halide perovskites, including RP structures, and hybrid compositions, when compared to HSE06. They further show the band gaps predicted by DeePKS and PBE, with respect to those by HSE06, over 30 tested systems. As depicted in Fig. 22, The DeePKS model demonstrated precise predictions of band gaps for all perovskites examined, closely aligning with the HSE06 findings and yielding an average absolute error (MAE) of 0.0350 eV. On the other hand, PBE significantly underestimated the band gap values, with a large MAE of 0.5222 eV.
@@
Zhang et al. Zhang et al. (2024b) investigated the tautomeric equilibria of glycine in water with the DeePMD model at the accuracy of M06-2XZhao and Truhlar (2008). To avoid expensive computational cost for directly generating dataset at M06-2X level, they utilized the DeePKS method as implemented in ABACUS and DeePKS-kit, together with on-the-fly probability enhanced sampling (OPES)Invernizzi and Parrinello (2020) method, to construct the dataset for training the DeePMD model. With this DeePMD potential and OPES, they performed MD simulations and obtained a converged free energy surface (FES). They observed that glycine can undergo tautomerism, transitioning between its neutral and zwitterionic forms through intramolecular and intermolecular proton transfers. In Zhang and Xu (2025), Zhang et al. adopted a similar strategy to study the propensity of water self-ions at air(oil)-water interface. They found that the trained DeePKS model can decrease calculation time by about nine times compared to regular M06-2X calculations. With the resulting DeePMD model for an efficient MD process, they demonstrated the stable ionic double-layer distribution near the interface for both air-water and oil-water interface systems.
@@ IV.6 DFT+U
where $E_{\text{LDA}}$ denotes the baseline energy. In addition, the Hubbard term and double-counting terms are labeled as $E_{\mathrm{Hub}}$ and $E_{\mathrm{dc}}$, respectively. In ABACUS, we introduce the fully localized limit (FLL) implementation of DFT+$U$ method Dudarev et al. (1998), which assumes that the on-site Coulomb interactions of localized electrons are fully accounted for by the Hubbard $U$ term, while a mean-field average is subtracted to avoid overestimating these interactions. The Hubbard-corrected energy functional is given by
@@
Table 1 presents the calculated band gaps and atomic magnetic moments for transition metal oxides (TMOs) with the Hubbard U parameter ranging from 0 to 6 eV. All of the four investigated TMOs (MnO, FeO, CoO, and NiO) are in the rhombohedral structure with a type-II antiferromagnetic (AFM) configuration. The lattice parameters, adopted from Ref. Tran et al., 2006, are 4.445, 4.334, 4.254, and 4.171 Å for MnO, FeO, CoO, and NiO, respectively, with the AFM ordering aligned along the [111] crystallographic direction. Fig. 25 further shows the comparison of band gaps for MnO, FeO, CoO, and NiO systems. The results obtained from ABACUS match well with those from other DFT packages such as VASP, Wien2k, and Quantum ESPRESSO Qu et al. (2022a).
@@ IV.7 Real-Time Time-Dependent DFT
Figure 26: Real-time TDDFT applications on photocatalysis Ren et al. (2024). (a) Projected density of states for H2O on Ni single atom site load on CeO2. (b) Bond length of thermal dissociation of H2O under 600 K. (c-d) Bond length of light-induced dissociation of H2O. (Adapted with permission from Nat. Commun. 15, 4675 (2024). Copyright 2024 Springer Nature.)
@@ V ABACUS in the OpenLAM Project
As the collection of quantum mechanical data progressively encompasses the entire periodic table, the Deep Potential team has launched an ambitious project named the Large Atomic Model (LAM) based on the practice of DPA pre-training model Zhang et al. (2024c, d). Emphasizing open scientific collaboration, this initiative has been formally designated as the OpenLAM project ope (2024a), reflecting its commitment to transparent development and community-driven progress. ABACUS serves as the cornerstone computational platform for generating cost-effective yet high-fidelity first-principles data to support OpenLAM’s development. Through heterogeneous algorithm optimization specifically tailored for plane-wave basis solutions of the Kohn-Sham equations, the software achieves remarkable efficiency gains in data generation.
@@ V.1 APNS Project
For those DFT codes that require pseudopotentials, several open-source codes for generating pseudopotentials are available, enabling users to tailor pseudopotentials to their specific requirements Schlipf and Gygi (2015); Holzwarth et al. (2022). For LCAO calculations using ABACUS, it is crucial to validate both the NAO basis set and the pseudopotentials before practical applications. To address the above issues, we have initiated the "ABACUS Pseudopotential Numerical Atomic Orbital Square" (APNS) project.
@@
Systematic verification of ABACUS remains essential in the APNS project to ensure reliable software development and maintain efficient high-throughput calculations. This aligns with growing research on validating DFT codes Talman (1978); Toyoda and Ozaki (2010); Wang et al. (2019b); Lejaeghere et al. (2016). Major databases like Materials Project Jain et al. (2020), Materials Cloud Talirz et al. (2020), and Computational Materials Repository Landis et al. (2012) have improved DFT practices. These datasets help balance accuracy and efficiency in DFT calculations while establishing code implementation standards. The APNS workflow enables automated high-throughput testing with optional user interaction. It evaluates pseudopotentials and NAO basis set across predefined and customizable systems, compatible with both PW and LCAO methods. The system also maintains a results database for test verification and validated pseudopotential/orbital file downloads. Here we list some examples done in the APNS project.
@@
Fig. 27 demonstrates the standardized testing workflow, where the workflow has conducted efficiency and accuracy evaluations for multiple norm-conserving pseudopotential families including SG15 SG15 ONCV Potentials , PseudoDojo Pseudo Dojo , PD03/PD04 PWmat-website (with version variations and semi-core configurations), CP2K’s Goedecker-Teter-Hutter pseudopotentials Krack (2005), and three ultrasoft types (pslibrary 0.3.1/1.0.0 theos-pp , GBRV v1.5 Garrity et al. (2014); GBRV pseudopotentials ). This workflow has additionally supported ABACUS feature development such as DFT+U implementation and ultrasoft pseudopotential compatibility.
@@
Figure 27: Workflow-chart of ABACUS Pseudopotential-Numerical atomic orbital Square automatic (APNS), high-throughput workflow incorporating structure generation, download and management (interfaced with online open crystal structure databases Materials Project Jain et al. (2020), Optimade Andersen et al. (2021), Crystallography Open Database (COD) Gražulis et al. (2009), etc.), symmetry analysis (interfaced with SeeK-path Togo, Shinohara, and Tanaka (2024); Hinuma et al. (2017)), modular and extensible software input generation (ABACUS and Quantum ESPRESSOGiannozzi et al. (2020)) and remote high-throughput computing support.
@@
The PW kinetic energy cutoff critically determines calculation accuracy, but excessively high values hinder computational efficiency. To address this trade-off, we established a systematic protocol for optimizing pseudopotentials through convergence testing across key parameters. This process involves two-phase validation. First, convergence testing of cutoff energies (20-300 Ry range) using stable crystal structures from Materials Project Talirz et al. (2020), Crystallography Open Database Gražulis et al. (2009), and Optimade Andersen et al. (2021). Second, equation of state (EOS) verification with optimized cutoffs. Three physical properties, including Kohn-Sham energy, lattice pressure (stress tensor trace), and band structure similarity metrics Prandini et al. (2018) were used as convergence criteria across pseudopotential types.
@@
EOS tests were performed using non-spin-polarized calculations, with $k$-sampled at a spacing of 0.06 Bohr-1 in the Brillouin zone for all elements. The crystal volumes $V$ range from 94% ($V_{m}$) to 106% ($V_{M}$) with a 2% step size. The kinetic energy cutoff for the plane-wave (PW) basis is set to the maximum of the converged values for all elements in the tested system. The $\Delta$ value Lejaeghere et al. (2014), defined as
@@
serves as a scalar indicator to quantify the difference between EOS profiles derived from pseudopotential and all-electron calculations (denoted by $a$ and $b$, respectively). The continuous integral in Eq. (113) is evaluated over the curve fitted using the Birch-Murnaghan equation Birch (1947), where $E_{a}(V)$ and $E_{b}(V)$ represent the Kohn-Sham energies relative to their equilibrium values.

### v4

@@ IV.3 Forces and Stress
where the kinetic and nonlocal pseudopotential matrices are respectively labeled as $T_{\mu\nu}(\mathbf{R})$ and $V_{\mu\nu}^{\rm NL}(\mathbf{R})$, while the local pseudopotential is $V_{\rm local}$ and the ionic energy $E_{\rm II}$ is given by the Ewald method. Martin (2020) It should be noted that ABACUS uses Eq. 33 to calculate the total energy in practical computations. For more information about evaluating force and stress terms in ABACUS, we refer the readers to a recent review work in Ref. Lin et al., 2024.
@@ IV.4 Hybrid Functional
Hybrid density functionals (HDFs) formulated within the generalized KS framework Seidl et al. (1996) belongs to the fourth rung of the Jacob’s ladder. Perdew and Schmidt (2001) It overcomes the drawbacks of the local and semi-local XC functionals, namely the self-interaction errors Perdew and Zunger (1981); Mori-Sánchez, Cohen, and Yang (2008) related to the underestimation of the band gaps in solids. The ABACUS package supports pure exact exchange (EXX) calculations, specifically the Hartree-Fock (HF) method. Building upon this capability, we have implemented several widely-used hybrid functionals, including the Heyd-Scuseria-Ernzerhof (HSE), Heyd, Scuseria, and Ernzerhof (2003) PBE0, Ernzerhof and Scuseria (1999); Adamo and Barone (1999b) and SCAN0. Hui and Chai (2016) The local (semilocal) components of these functionals are computed using the LibXC library, Lehtola et al. (2018) while the exact (Fock) exchange energy is evaluated through two-electron integrals of the Kohn-Sham orbitals.
@@
In fact, the above EXX effective potential exhibits an unfavorable $O(N^{4})$ scaling behavior, making it significantly more computationally demanding than local functionals in terms of both time and memory requirements. To mitigate this challenge, several algorithmic approaches have been developed, most notably the density fitting (DF) technique Whitten (1973) and its specialized implementation for localized atomic orbitals known as resolution of identity (RI). Feyereisen, Fitzgerald, and Komornicki (1993); Vahtras, Almlöf, and Feyereisen (1993); Ren et al. (2012) These methods reduce the four-center integrals to combinations of three- and two-center integrals by expanding the orbital products using auxiliary basis functions (ABFs).
For HDF calculations in large periodic systems, the local resolution of identity (LRI) approach Levchenko et al. (2015); Ihrig et al. (2015); Lin, Ren, and He (2020) enables linear-scaling EXX calculations by exploiting the inherent locality of both atomic orbitals (AOs) and ABFs. Extensive benchmarking has demonstrated that with properly chosen ABFs, the LRI approximation achieves sufficient accuracy for HDF calculations. Lin, Ren, and He (2020) Residual discrepancies in cohesive properties and band gaps primarily originate from differences in core-valence interaction treatments (pseudopotentials versus all-electron methods) rather than basis set limitations. Ji et al. (2022)
@@
Figure 19: Band gaps of 21 different semiconductors, calculated by ABACUS+LibRI, along with the reference computed values from other software (QE/FHI-AIMS/VASP) Lin, Ren, and He (2020) and experimental values. Cui et al. (2018)
The computational efficiency is substantially enhanced by exploiting the inherent locality of both AOs and the density matrix, which enables effective prescreening of negligible submatrices below a specified threshold. For parallel computation of four-center integrals, we have implemented two complementary distribution strategies: a load-balancing greedy algorithm for multi-processor scheduling, and a spatially optimized scheme based on K-means clustering of ABF atom pairs. The latter minimizes both inter-processor communication overhead and memory requirements by ensuring geometrically proximate pairs are assigned to the same processor, with memory consumption scaling linearly with the union of their neighbor lists. Lin, Ren, and He (2021a)
Notably, the LRI-based four-center integration has been modularized into the standalone LibRI library, lib (2024a) which provides a portable and efficient implementation that can be readily integrated into various electronic structure codes. This includes both conventional DFT packages and beyond-DFT methods such as the random phase approximation (RPA) and GW calculations, lib (2024b) offering substantial acceleration for two-electron Coulomb repulsion integral evaluations across different theoretical frameworks.
In 2025, Lin et al. developed a linear-scaling algorithm for exact-exchange forces and stresses in hybrid DFT, applicable to both molecules and periodic systems. Leveraging numerical atomic orbitals and localized resolution-of-identity, the massively parallel implementation enables structural relaxation of thousand-atom systems in ABACUS. Lin et al. (2025)
@@
Generally speaking, hybrid functional calculations can only process systems of limited size due to the large computational demand of building the exact exchange Hamiltonian. By combining NAOs basis and LRI techniques, Fig. 19 shows the time consumption of this part increases almost linearly with the system size, enabling ABACUS to perform hybrid functional calculations for large systems with thousands of atoms. In 2024, Tang et al. used hybrid functional data from ABACUS and trained a deep equivariant neural network approach for efficient hybrid density functional calculations. Tang et al. (2024) Taking twisted bilayer graphene (TBG) as an example in Fig. 20, they performed the HSE band structure calculations of (17, 16) TBG (twist angle $\theta\approx 2.0046^{\circ}$, 3,268 atoms/cell). Tang et al. (2024)
We also performed HSE calculations to obtain the band gaps of several semiconductors and compare them with the results from other DFT packages (QE/VASP/FHI-AIMS) and the experimental values. Lin, Ren, and He (2020); Cui et al. (2018) In these tests, ABACUS and QE employ the SG15-type norm-conserving PBE pseudopotentials (except for In is PSlibrary norm-conserving), VASP employs the PBE PAW potentials, whereas FHI-AMIS performs all-electron calculations. Delley (1990) Fig. 19 shows that the band gaps obtained by ABACUS HSE calculations are consistent with the results from other DFT packages, demonstrating satisfactory results when compared to the experimental values.
@@ IV.5 DeePKS
The exchange-correlation functional is critical to determining the accuracy of DFT. However, when different levels of XC functionals are selected, there exists a trade-off between accuracy and efficiency. Lower-level XC functionals on Jacob’s ladder Perdew (2001) typically offer higher computational efficiency but lower accuracy, whereas higher-level ones exhibit the opposite trend. With the rapid progress in increasing computational resources and advanced algorithms, AI-assisted methods hold the potential to mitigate this issue. Kulik et al. (2022); Huang, Von Rudorff, and Von Lilienfeld (2023)
Proposed in 2020, DeeP Kohn-Sham method (DeePKS)Chen et al. (2021) employs a computationally efficient neural network-based functional model to represent the energy and force difference between a lower-level XC functional and a higher-level XC functional. The resulting model maintains translational, rotational, and permutational invariance and can be used in self-consistent field calculations. Together with DeePKS-kit Chen et al. (2022a) software, ABACUS supports the iterative training of DeePKS models. Li et al. (2022a) Specifically, DeePKS models can be trained in molecular or periodic systems to learn properties such as energy, force, stress, and band gap. Based on the DeePKS model, ABACUS can achieve accuracy similar to that of the high-level XC functional on these specified properties, while the computational efficiency is similar to that of the low-level XC functional.
@@
Here we list a few applications using the DeePKS method. For instance, halide perovskites (ABX3, X=halogen anion) have shown great promise as a cost-effective alternative to current commercial photovoltaic technologies. Designing effective photovoltaic systems requires a precise yet efficient description of the electronic structure of halide perovskites. One key benefit of halide perovskites is adjusting the absorption edge wavelength (band gap) by changing the ratio of different halide ions. Ou et al. Ou et al. (2023) constructed a general DeePKS model that can be utilized for halide perovskites, including different combinations of ABX3 (A=FA, MA, Cs; B=Sn, Pb; X=Cl, Br, I), the organic-inorganic hybrid alternatives, and the Ruddlesden-Popper (RP) perovskites. In detail, the authors built an extensive DeePKS model upon 460 configurations spanning seven types of halide perovskites, with HSE06 accuracy and satisfactory predictions for the band gap. Based on an iterative training process with ABACUS and DeePKS-kit, they showed that the resulting DeePKS model can accurately replicate forces, stress, band gaps, and density of states (DOS) near the Fermi energy for all types of halide perovskites, including RP structures, and hybrid compositions, when compared to HSE06. They further showed the band gaps predicted by DeePKS and PBE, with respect to those by HSE06, over 30 tested systems. As depicted in Fig. 21, the DeePKS model demonstrated precise predictions of band gaps for all perovskites examined, closely aligning with the HSE06 findings and yielding an average absolute error (MAE) of 0.0350 eV. On the other hand, PBE significantly underestimated the band gap values, with a large MAE of 0.5222 eV.
@@
Zhang et al. Zhang et al. (2024c) investigated the tautomeric equilibria of glycine in water with the DeePMD model at the accuracy of M06-2X. Zhao and Truhlar (2008) To avoid expensive computational cost for directly generating dataset at M06-2X level, they utilized the DeePKS method as implemented in ABACUS and DeePKS-kit, together with on-the-fly probability enhanced sampling (OPES) Invernizzi and Parrinello (2020) method, to construct the dataset for training the DeePMD model. With this DeePMD potential and OPES, they performed MD simulations and obtained a converged free energy surface (FES). They observed that glycine can undergo tautomerism, transitioning between its neutral and zwitterionic forms through intramolecular and intermolecular proton transfers. Zhang et al. Zhang and Xu (2025) adopted a similar strategy to study the propensity of water self-ions at air(oil)-water interface. They found that the trained DeePKS model can decrease calculation time by about nine times compared to regular M06-2X calculations. With the resulting DeePMD model for an efficient MD process, they demonstrated the stable ionic double-layer distribution near the interface for both air-water and oil-water interface systems.
@@ IV.6 DFT+U
where $E_{\text{DFA}}$ denotes the baseline energy. In addition, the Hubbard term and double-counting terms are labeled as $E_{\mathrm{Hub}}$ and $E_{\mathrm{dc}}$, respectively. In ABACUS, we introduce the fully localized limit (FLL) implementation of DFT+$U$ method, Dudarev et al. (1998) which assumes that the on-site Coulomb interactions of localized electrons are fully accounted for by the Hubbard $U$ term, while a mean-field average is subtracted to avoid overestimating these interactions. The Hubbard-corrected energy functional is given by
@@
Table 1 presents the calculated band gaps and atomic magnetic moments for transition metal oxides (TMOs) with the Hubbard U parameter ranging from 0 to 6 eV. All of the four investigated TMOs (MnO, FeO, CoO, and NiO) are in the rhombohedral structure with a type-II antiferromagnetic (AFM) configuration. The lattice parameters, adopted from Ref. Tran et al., 2006, are 4.445, 4.334, 4.254, and 4.171 Å for MnO, FeO, CoO, and NiO, respectively, with the AFM ordering aligned along the [111] crystallographic direction. Fig. 24 further shows the comparison of band gaps for MnO, FeO, CoO, and NiO systems. The results obtained from ABACUS match well with those from other DFT packages such as VASP, Wien2k, and Quantum ESPRESSO. Qu et al. (2022a)
@@ IV.7 Real-Time Time-Dependent DFT
Figure 25: Real-time TDDFT applications on photocatalysis. Ren et al. (2024) (a) Projected density of states for H2O on Ni single atom site load on CeO2. (b) Bond length of thermal dissociation of H2O under 600 K. (c-d) Bond length of light-induced dissociation of H2O. (Adapted with permission from Nat. Commun. 15, 4675 (2024). Copyright 2024 Springer Nature.)
@@ V ABACUS in the OpenLAM Project
As the collection of quantum mechanical data progressively encompasses the entire periodic table, the Deep Potential team has launched an ambitious project named the Large Atomic Model (LAM) based on the practice of DPA pre-training model. Zhang et al. (2024d, e) Emphasizing open scientific collaboration, this initiative has been formally designated as the OpenLAM project, ope (2024a) reflecting its commitment to transparent development and community-driven progress. ABACUS serves as the cornerstone computational platform for generating cost-effective yet high-fidelity first-principles data to support OpenLAM’s development. Through heterogeneous algorithm optimization specifically tailored for plane-wave basis solutions of the Kohn-Sham equations, the software achieves remarkable efficiency gains in data generation.
@@ V.1 APNS Project
For those DFT codes that require pseudopotentials, several open-source codes for generating pseudopotentials are available, enabling users to tailor pseudopotentials to their specific requirements. Schlipf and Gygi (2015); Holzwarth et al. (2022) For LCAO calculations using ABACUS, it is crucial to validate both the NAO basis set and the pseudopotentials before practical applications. To address the above issues, we have initiated the "ABACUS Pseudopotential Numerical Atomic Orbital Square" (APNS) project.
@@
Systematic verification of ABACUS remains essential in the APNS project to ensure reliable software development and maintain efficient high-throughput calculations. This aligns with growing research on validating DFT codes. Talman (1978); Toyoda and Ozaki (2010); Wang et al. (2019b); Lejaeghere et al. (2016) Major databases like Materials Project, Jain et al. (2020) Materials Cloud, Talirz et al. (2020) and Computational Materials Repository Landis et al. (2012) have improved DFT practices. These datasets help balance accuracy and efficiency in DFT calculations while establishing code implementation standards. The APNS workflow enables automated high-throughput testing with optional user interaction. It evaluates pseudopotentials and NAO basis set across predefined and customizable systems, compatible with both PW and LCAO methods. The system also maintains a results database for test verification and validated pseudopotential/orbital file downloads. Here we list some examples done in the APNS project.
@@
Fig. 26 demonstrates the standardized testing workflow, where the workflow has conducted efficiency and accuracy evaluations for multiple norm-conserving pseudopotential families including SG15, SG15 ONCV Potentials PseudoDojo, Pseudo Dojo PD03/PD04 PWmat-website (with version variations and semi-core configurations), CP2K’s Goedecker-Teter-Hutter pseudopotentials, Krack (2005) and three ultrasoft types (pslibrary 0.3.1/1.0.0, theos-pp GBRV v1.5 Garrity et al. (2014); GBRV pseudopotentials ). This workflow has additionally supported ABACUS feature development such as DFT+U implementation and ultrasoft pseudopotential compatibility.
@@
Figure 26: Workflow-chart of ABACUS Pseudopotential-Numerical atomic orbital Square automatic (APNS), high-throughput workflow incorporating structure generation, download and management (interfaced with online open crystal structure databases Materials Project, Jain et al. (2020) Optimade, Andersen et al. (2021) Crystallography Open Database (COD), Gražulis et al. (2009) etc.), symmetry analysis (interfaced with SeeK-path Togo, Shinohara, and Tanaka (2024); Hinuma et al. (2017)), modular and extensible software input generation (ABACUS and Quantum ESPRESSO Giannozzi et al. (2020)) and remote high-throughput computing support.
@@
The PW kinetic energy cutoff critically determines calculation accuracy, but excessively high values hinder computational efficiency. To address this trade-off, we established a systematic protocol for optimizing pseudopotentials through convergence testing across key parameters. This process involves two-phase validation. First, convergence testing of cutoff energies (20-300 Ry range) using stable crystal structures from Materials Project, Talirz et al. (2020) Crystallography Open Database, Gražulis et al. (2009) and Optimade. Andersen et al. (2021) Second, equation of state (EOS) verification with optimized cutoffs. Three physical properties, including Kohn-Sham energy, lattice pressure (stress tensor trace), and band structure similarity metrics Prandini et al. (2018) were used as convergence criteria across pseudopotential types.
@@
EOS tests were performed using non-spin-polarized calculations, with $k$-sampled at a spacing of 0.06 Bohr-1 in the Brillouin zone for all elements. The crystal volumes $V$ range from 94% ($V_{m}$) to 106% ($V_{M}$) with a 2% step size. The kinetic energy cutoff for the plane-wave (PW) basis is set to the maximum of the converged values for all elements in the tested system. The $\Delta$ value, Lejaeghere et al. (2014) defined as
@@
serves as a scalar indicator to quantify the difference between EOS profiles derived from pseudopotential and all-electron calculations (denoted by $a$ and $b$, respectively). The continuous integral in Eq. (113) is evaluated over the curve fitted using the Birch-Murnaghan equation, Birch (1947) where $E_{a}(V)$ and $E_{b}(V)$ represent the Kohn-Sham energies relative to their equilibrium values.